OpenStudy (anonymous):

Solve the equation in the intervals [0,2pi) 4sin^2x+2cos^2x=3

OpenStudy (anonymous):

\[4\sin^2x+2\cos^2x=3\]

OpenStudy (fools101):

I here try this web site should work! http://www.mathway.com Let me know if this don't help you! ~Jess~

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it didnt give the exact values

OpenStudy (fools101):

oh okay... Sorry

OpenStudy (anonymous):

Thanks anyway

OpenStudy (fools101):

ya i am sorry i won't that good as a help, let me look for other on okay! = )

OpenStudy (fools101):

here try this web site! http://www.wolframalpha.com/